# Disentangling Representations in RNNs through Multi-task Learning

### v1

@@ 4 Experiments
4.1 Multitasking leads to disentangled representations
We train RNNs to do simultaneous classifications for 24 linear partitions of the latent space for $D=2$ (Figure 1b, 6 partitions shown). Figure 1d shows the top 3 PCs (capturing $\sim 85\,\%$ of the variance) of network activity after training (final accuracies $\sim 95\%$) for multiple trials, along with the fixed points of network dynamics. To find the fixed points, we follow a standard procedure outlined in Sussillo and Barak [69] (see Section A.1 for details). Looking at Figure 1d the fixed points span the entire two-dimensional manifold that the trials evolve in, which corresponds to a continuous attractor with stable states across a 2D "sheet". This means that the network can store a short-term memory [73] of the current amount of accumulated evidence in a product space of the latent variables.
Figure 1: Learning disentangled representations. (a) A disentangled representation directly lends itself to OOD generalization: a downstream linear decoder that can differentiate ripe from unripe bananas can readily generalize to mangos, even though it has never been trained on mangos. (b) The task is to simultaneously report whether the ground truth $\mathbf{x}^{*}$ lies above ($+1$) or below ($-1$) a number of classification lines. (c) RNNs are trained to report the outcome of all the binary classifications in b at the end of the trial (indicated by the fixation input turning 0). (d) Top 3 PCs of RNN activity. Each line is a trial, while color saturation indicates time. All trials start from the center and move outwards, towards the location of $\mathbf{x}^{*}$ in state space. We color the last timepoint in each trial (squares) according to the quadrant this trial was drawn from. Red x’s correspond to attractors (see Section A.1). Here we remove input noise so that trajectories can be visualized easier. The network learns a two-dimensional continuous attractor that seems to provide a disentangled representation of the state space. (e) To evaluate OOD generalization, we train a linear decoder (see a) to output the ground truth $\mathbf{x}^{*}$ at the end of the trial, while keeping network weights frozen. (f) ID and OOD generalization performance for networks trained in different number of tasks $N_{\textit{task}}$. We report the 25, 50 and 75 percentile of $r^{2}$ for each network size (see Section A.2). ID and OOD performance increase with number of tasks, and the generalization gap decreases, indicating that the networks have indeed learned abstract representations. (g) The results hold for other recurrent architectures.
Furthermore, the representation is low-dimensional, rectangular and compositional; all characteristics of disentangled representations. In comparison, the representations after the encoder are non-linearly mixed, high-dimensional and overlapping (Figure S5a). Individual trials with noise show how the representation maintains a sense of metric distances in the RNN representation space (Figure S5b). Figure S5c demonstrates how this representation comes about during learning, and Figure S5d that the short-term memory persists when a delay period is included before the decision. Therefore, it seems that multitasking has led to disentagled, persistent representations of the latent variables. Importantly, and in line with our theory, this only happens when noise is present in the input, which forces the network to learn a notion of distance from classification boundaries. Indeed, when the network is trained without input noise, it does not learn a 2D continuous attractor (Figure S5e).
@@ 4.1 Multitasking leads to disentangled representations
To quantify the disentanglement of the representations, we evaluate regression generalization by training a linear decoder to predict the ground truth $\mathbf{x}^{*}$ while network weights are frozen (Figure 1e). We perform out-of-distribution 4-fold crossvalidation, i.e. train the decoder on 3 out of 4 quadrants and test in the remaining quadrant (Section A.2 for details). We also evaluate in-distribution (ID) performance by training the decoder in all quadrants. An example of train and test losses is shown in Figure S5f. We find that the network’s OOD and ID generalization performance are excellent (median $r^{2}=0.96,0.97$ respectively across 5 example networks, also see Figure 1f); therefore the network has indeed learned a disentangled representation that zero-shot generalizes OOD. In addition, ID performance increases with the number of tasks $N_{\textit{task}}$, and the OOD generalization gap decreases (Figure 1f). Therefore we conclude that simultaneously training an RNN in classification tasks that span the latent space results in disentangled representations in the recurrent hidden layer.
@@
Since $\mathbf{x}^{*}$ can be decoded by this representation in unseen (by the decoder) parts of the state space, it follows that the representation can be used to solve any task involving the same underlying variables, without requiring further pretraining. In other words, to solve any other task we do not need to deal with the denoising and unmixing of the latent factors $x_{1}$, $x_{2}$; we would just need to learn the (non-linear) mapping from $x_{1}$, $x_{2}$ to task output. Crucially, these findings are architecture-agnostic: they hold for non-leaky ("vanilla") RNNs, which outperform leaky ones, and LSTMs which drastically reduce performance variability (Figure 1g). We still choose leaky RNNs because of their closer correspondence to biological neurons, which have a membrane voltage that decays over time.
We were also curious to see the impact of correlated inputs. A problem with high correlations is that they render parts of the state space virtually invisible to the network (Figure S6a). Surprisingly, OOD generalization performance is very weakly affected by input correlations, even though the state space is sampled uniformly in test (Figure S6b). The behavior is highly non-linear: performance is great until $\rho=0.97$, but when the inputs are perfectly correlated ($\rho=1$), performance drop is sharp.
@@
Finally, we examined RNN unit activations for these representations, and their relation to the latent variables. In Figure S7a we plot the steady-state firing rates for all 64 neurons, while regularly sampling $x_{1}$ and $x_{2}$. Interestingly, we find that in these networks only $\sim 10\,\%$ of neurons are active at any time, which is in line with sparse coding in the brain. In addition, the average firing rate is consistently $\sim 0.25$ spikes/s, which is surprisingly close to cortical values. Furthermore, we find that all of the active neurons display mixed selectivity, i.e. they are tuned to both variables, which is a known property of cortical neurons [61] (Figure S7b). This suggests that metrics of disentanglement that assume that individual neurons encode distinct factors of variation [29, 39, 13, 32, 21] might be insufficient in detecting disentanglement in networks that generalize well. While recent work incorporates such axis-alignment in the definition of disentaglement, our work along with others [38] showcases the advantages of approaching disentanglement from a mixed representations perspective. Importantly, these properties were not imposed during training, nor was there any parameter fine tuning involved; rather they emerged from task structure and optimization objective.
4.2 Nonlinear tasks and interleaved learning
So far we have been training RNNs to perform linear classifications. However there are cases where the streams might need to be combined nonlinearly. For instance, if the two streams represent the amount and probability of reward respectively, an agent needs to multiply the two and decide whether the expected value exceeds a certain (metabolic) cost $\gamma$ of performing an action to obtain said reward. Figure S8a shows the classification lines for the multiplicative task, where the network should decide whether the ground truth $\mathbf{x}^{*}$ lies above or below the curve $x_{1}\,x_{2}=\gamma$, for multiple values of $\gamma$. RNNs trained on this task learn a representation that separates for trials in different quadrants (Figure S8b). This heuristic might have come about because individual classification lines span only a single quadrant for this task. Furthermore, trials belonging to the same quadrant seem to be more convoluted, with a corresponding drop in OOD performance (median $r^{2}=0.71$), which might reflect the non-linearity of the task. Therefore, we wondered how the representation would look like if the network was trained on both the linear and multiplicative tasks, as animals need to do.
For that we perform interleaved training of both tasks (i.e. train in batches sampled from one of the tasks at a time), a setting where neural networks excel at, compared to humans who excel at blocked training, where tasks are learned sequentially (but see Flesch et al. [24]). Figure S8c shows that the network learns a two-dimensional continuous attractor, and the learned representation is no longer separated by quadrants. OOD generalization for this network is excellent, and almost identical to ID performance (median $r^{2}=0.94,0.97$ respectively). Overall, we conclude that our framework extends to interleaved learning of a mixture of linear and nonlinear tasks, which better reflects the challenges encountered by agents in the real world. Note that during interleaved training, linear and non-linear tasks are not performed simultaneously; yet they are in immediate succession which also places pressure to the network to gradually learn representations that satisfy all tasks. The relationship between interleaved learning and multitasking could be an interesting direction for future work.
4.3 Abstract representations are learned for a free reaction time, integrate to bound task
Figure 2: Free reaction time task. (a) Data generating process. Every classification line from Figure 1b now corresponds to an accumulator (see corresponding colors), and the desired output for the RNN is the accumulator values for the entire trial. The accumulator is quantized to integer values between ±$\pm 5$. (b) Representation for RNN trained on free reaction time task. The network learns a two-dimensional continuous attractor, similar to Figure 1d. (c) OOD generalization performance for the free reaction time (RT) task. Free RT outperforms fixed RT for a small number of tasks.
So far we have been training networks to produce a response at the end of the trial. However, in many situations agents are free to make a decision whenever they are certain enough. Therefore, we here seek to extend our framework to free reaction time (RT) decisions. A canonical model accounting for choices and reaction times in humans and animals is the drift-diffusion model [41, 11]. It is composed of an accumulator that integrates noisy evidence over time, until a certain amount of certainty, represented by a bound, is reached, triggering a decision. In the linear classification task setting, the accumulated amount of evidence at time $t$ for a line with slope $\alpha$, $A_{\alpha}(t)$ is given by:
$A_{\alpha}(t)=A_{\alpha}(t-1)+X_{1}(t)-\alpha\,X_{2}(t)$ (5)
Intuitively, $A_{\alpha}(t)$ reflects the amount of confidence at time $t$ that the ground truth $\mathbf{x}^{*}$ lies above or below the classification line with slope $\alpha$. Essentially, the network has to explicitly report distance from the classification lines, not just in which side of the line $\mathbf{x}^{*}$ lies for that trial. We arbitrarily set the decision bound to ±$\pm 5$, and plot the accumulators $A_{\alpha}$ for all lines in Figure 2a. Note that once the bound is reached a decision is effectively made and $A_{\alpha}$ is kept constant. Also, instead of using continuous values, we quantize $A_{\alpha}$, because it is going to be used as target signal to train the network, and we do not want to introduce a strong inductive bias towards integrating the evidence streams.
We then train the RNN to reproduce confidence estimates from Figure 2a for the entire trial. Compared to previous experiments, the fixation input is no longer available to determine when to produce a decision. Instead, decisions evolve dynamically throughout the trial. Figure 2b shows that in this setting the network still learns a two-dimensional continuous attractor of the latent space. Furthermore, the free RT outperforms the fixed RT network (Figure 2c) for a small number of tasks, since it is explicitly required to report distance from the classification lines. However, as our theory shows (Lemma B.3) the fixed RT network is also implicitly reporting distance from the boundaries, when behaving like an optimal multi-task classifier, which explains the similar performance for a larger number of tasks. Overall, we showed that our setting accounts for naturalistic free RT decisions, and provides theoretical justification for the importance of confidence signals in the brain [64, 49].
4.4 RNNs confirm and extend theoretical predictions
@@
Here we expore the relation between the theory in Section 3 and Appendix B and experiments in Section 4 in more depth. First, we wondered why performance saturates in our networks to a high yet non-1 $r^{2}$. The central limit theorem predicts that the estimate of the ground truth $\mathbf{x}^{*}$ in any optimal multi-task classifier becomes more accurate with $\sqrt{t}$, providing a theoretical maximum $r^{2}$ given trial duration $T$ (Section A.3). Since the networks trained on the free RT task are required to output their confidence at any time in the trial, we can compute OOD $r^{2}$ of network predictions at any timepoint $t$, and compare that to the theoretical prediction. Figure S9a shows that indeed the highest RNN $r^{2}$ falls in the vicinity of or just short of the theoretical maximum. This indicates that RNNs trained with BPTT on these tasks behave like near-optimal multi-task classifiers that create increasingly accurate predictions with time, tightening the relation between our theoretical and experimental results.
An important prediction of our theory is that to learn abstract representations $N_{\textit{task}}$ should exceed $D$. To test this, we increase $D$ (adding more inputs to Figure 1c), and vary $N_{\textit{task}}$. Sampling classification hyperplanes methodically in high-dimensional spaces is non-trivial; therefore we resort to randomly sampling them. Figure 3a shows OOD generalization performance for various combinations of $D$ and $N_{\textit{task}}$. We observe that performance is bad when the $N_{\textit{task}}<D$, but it increases when $N_{\textit{task}}\geq D$. This increase is abrupt for smaller $D$ and more gradual for higher, which is in line with remarks by Johnston and Fusi [38] that it is easier to learn abstract representations when $D$ is high. Overall, our findings confirm our theory that abstract representations emerge when $N_{\textit{task}}\geq D$. This result is remarkable, especially for high $D$, because it goes against our intuition that $N_{\textit{task}}$ should scale exponentially with $D$ to fill up the space adequately; instead it need only scale linearly.
@@
Figure 3: RNNs confirm theoretical predictions. (a) OOD $r^{2}$ as a function of input dimensionality $D$ and number of tasks $N_{\textit{task}}$. Good values of $r^{2}$ are obtained when $N_{\textit{task}}\geq D$, confirming our theoretical results. (b) Increasing amounts of noise in pretraining results in better OOD generalization.
@@
Our theory and experiments provide insight on the importance of noise for developing efficient, abstract representations (Figure S5e). The closer to a classification boundary the ground truth $\mathbf{x}^{*}$ is, the more likely noise will cross over the boundary. Since, as our theory shows, any optimal multi-task classifier has to estimate $\Pr\{y_{i}(\mathbf{x}^{*})=+1\}$, and said probability directly relates to the actual distance from the boundary, it follows that noise allows the model to learn distances from boundaries, leading to efficient localization. We reasoned that additional noise might be even more beneficial, as it would allow more accurate estimation of $\Pr\{y_{i}(\mathbf{x}^{*})=+1\}$, especially when $\mathbf{x}^{*}$ is far from the boundary. To test this, we increase noise strenght $\sigma$ during pretraining, while testing with the same $\sigma=0.2$. Indeed, increasing amounts of noise consistently result in better OOD generalization (Figure 3b). This benefit comes for smaller numbers of tasks, allowing us to consider less supervised tasks (e.g. 3), train on them with more noise, and achieve the same performance as more tasks (e.g. 12). So even though networks with more noise perform worse in pretraining (low 90%s classification accuracy), they learn more disentangled representations. These findings are highly non-trivial, and have informed our thinking about generalization and inherent variability of the underlying latent factors.
Finally, we were curious to see what happens if part of the state space was not crossed by task boundaries. To test this, we train a network for $D=2$ on classification tasks that only tile quadrants 1 and 3 (Figure S9b). Note that trials are still drawn from the entire state space – however the network need not differentiate between points drawn within quadrant 2 or 4. The network still learns a two-dimensional continuous attractor; however the representation is no longer disentangled, likely because the boundaries are far away if $\mathbf{x}^{*}$ is in quadrants 2 or 4 (Figure S9c). A more entangled representation may lead to poorer generalization performance, as measured by OOD regression generalization [38]. Indeed, we find that OOD performance of the tangled network is considerably worse (median $r^{2}=0.80$), and most importantly it performs badly in the quadrants that are not spanned by classification lines (25 percentile $r^{2}=0.62$). Therefore, it appears that in order to prevent the representations from tangling, tasks boundaries need to be distributed over the state space – as previously shown randomly sampling these boundaries is adequate.

### v2

@@ 5 Experiments
5.1 Multi-task learning leads to disentangled representations
@@
We train RNNs to do simultaneous classifications for $N_{\text{task}}$ linear partitions of the latent space for $D=2$ (Figure 2a, 6 partitions shown). To quantify the disentanglement of the representations after learning, we evaluate regression generalization by training a linear decoder to predict the ground truth $\mathbf{x}^{*}$ while network weights are frozen. We perform out-of-distribution 4-fold crossvalidation, i.e. train the decoder on 3 out of 4 quadrants and test in the remaining quadrant (Section A.2 for details). We also evaluate in-distribution (ID) performance by training the decoder in all quadrants. An example of train and test losses is shown in Figure S9f. We find that the network’s OOD and ID generalization performance are excellent (median $r^{2}=0.96,0.97$ respectively across 5 example networks); therefore the network has learned an abstract representation that zero-shot generalizes OOD. In addition, ID performance increases with the number of tasks $N_{\textit{task}}$, and the OOD generalization gap decreases (Figure 3a). Therefore we conclude that multi-task learning leads to abstract representations in the RNN’s hidden layer, when tasks span the latent space.
Figure 3: Learning disentangled representations. (a) ID and OOD generalization performance for networks trained in different number of tasks $N_{\textit{task}}$. We report the 25, 50 and 75 percentile of $r^{2}$ for each network size (see Section A.2). ID and OOD performance increase with $N_{\textit{task}}$, and the generalization gap decreases, indicating that the networks have learned abstract representations. (b) The results hold for other autoregressive architectures, including LSTMs and GPT-2 transformers. (c) Angles between latent factor decoders (see Section A.3 for how they were estimated). The angles approach 90 degrees as $N_{\text{task}}\gg D$ for RNNs, but already fror $N_{\text{task}}\geq D$ for transformers. Remaining errors around 90 degrees are attributed to variability in the linear decoder fits. (d) Top 3 PCs of RNN activity ($N_{\textit{task}}=24$, $D=2$). Each line is a trial, while color saturation indicates time. All trials start from the center and move outwards, towards the location of $\mathbf{x}^{*}$ in state space. We color the last timepoint in each trial (squares) according to the quadrant this trial was drawn from. Red x’s correspond to attractors (see Section A.6). Here we remove input noise so that trajectories can be visualized easier. The network learns a two-dimensional continuous attractor that provides a disentangled representation of the 2D state space. (e) Spectral plot resulting from linearizing RNN dynamics around every fixed point (Section A.6). First two eigenvalues of the difference system are near 00, while the rest decay much faster, indicating marginal stability across two dimensions for every fixed point, a signature of a 2D continuous attractor.
@@
Since $\mathbf{x}^{*}$ can be decoded by this representation in unseen (by the decoder) parts of the state space, it follows that the representation can be used to solve any task involving the same latent variables, without requiring further pretraining. In other words, to solve any other task we do not need to deal with the denoising and unmixing of the latent factors $x_{1}$, $x_{2}$; we would just need to learn the (potentially non-linear) mapping from $x_{1}$, $x_{2}$ to task output. Furthermore, the representation scales linearly with input dimensionality $D$ (see Figure 5b). This marks a significant improvement from previously proposed models for representation learning in the brain where one task is executed at a time [49, 78], which scale linearly with $N_{\textit{task}}$, and exponentially with $D$ (see Section A.7 for details). Crucially, these findings are architecture-agnostic: they hold for non-leaky ("vanilla") RNNs, which outperform leaky ones for small $N_{\textit{task}}$, LSTMs which perform the best, and GPT-2 transformers (details in Section A.5) which have excellent performance already from $N_{\textit{task}}=2$ (Figure 3b). Note that state-space models have superior asymptotic performance, which is expected due to the nature of the task. We focus on leaky RNNs because of their closer correspondence to biological neurons, which have a membrane voltage that decays over time.
So far we showcased abstractness, but not disentanglement. For disentaglement, it is crucial that the latents lie in orthogonal subspaces. Looking at the angles between the decoders of the latents, we find that they become orthogonal as $N_{\text{task}}\gg D$ for RNNs (Figure 3c), as predicted by our theory. Intriguingly, this already occurs from $N_{\text{task}}\geq D$ in transformers, showcasing their superior ability to separate latent factors. Furthermore, orthogonality strongly correlates with OOD generalization performance, which emphasizes the close link between disentanglement and abstractness: the more orthogonal the representations are, the cleaner the readout of the latent factors by linear decoders.
We further demonstrate the robustness of our setting by showing that abstract representations emerge for different noise distributions and correlated inputs (Section A.8), non-linear boundaries (Section A.9), and for cognitive neuroscience integrate-to-bound tasks where the agent can make their decision whenever confident enough, not at a fixed time [40] (Section A.10).
5.2 Representational structure in RNNs and Transformers
In this section, we open the black box and investigate the representations learned by the networks, starting with RNNs. Figure 3d shows the top 3 PCs (capturing $\sim 85\,\%$ of the variance) of network activity after training (final accuracy $\sim 95\%$) for multiple trials, along with the fixed points of network dynamics. To find the fixed points, we follow a standard procedure outlined in Sussillo and Barak [70] (see Section A.6 for details). Looking at Figure 3d the fixed points span the entire two-dimensional manifold that the trials evolve in, which corresponds to a continuous attractor with stable states across a 2D "sheet". Linearizing the dynamics around each fixed point and computing the eigenvalues of the linearized system (Section A.6 for details), reveals marginal stability across two eigenvectors, i.e. near-00 eigenvalues which correspond to slow, integration dimensions in network dynamics, therefore confirming the continuousness of the attractor (Figure 3e). This implies that the network can store a short-term memory [75] of the current amount of accumulated evidence in a product space of the latent variables, and update it as further evidence arrives.
Furthermore, compared to the representations after the encoder which are non-linearly mixed, high-dimensional and overlapping (Figure S9a), the representation in Figure 3d looks disentangled as we would expect from the theory and metrics above. Individual trials with noise show how the representation maintains a sense of metric distances in the RNN representation space (Figure S9b). Figure S9c demonstrates how this representation comes about during learning, and Figure S9d that the short-term memory persists when a delay period is included before the decision. Therefore, multi-task learning has led to disentangled, persistent representations of the latent variables. Importantly, and in line with our theory, this only happens when noise is present in the input, which forces the network to learn a notion of distance from classification boundaries (Lemma B.3). Indeed, when the network is trained without input noise, it does not learn a 2D continuous attractor (Figure S9e).
Figure 4: RNN and GPT representations and relation to latent variables. (a) Hidden layer activations of RNN in Figure 2b (left) and GPT-2 transformer (right), while systematically varying the latent factors $x_{1}$ and $x_{2}$ from -0.5 to 0.5. Activations are plotted in 8*8 grids, one for each value of $x_{1}$ and $x_{2}$. Each grid contains firing rates for a total of 64 neurons for the RNN, and activations for 8 neurons for each of the 8 heads from the final embedding of the sequence for GPT-2. (b) Correlation coefficient of activations for both models with $x_{1}$ and $x_{2}$, respectively.
@@ 5.2 Representational structure in RNNs and Transformers
Finally, we examined RNN and GPT unit activations, and their relation to the latent variables. In Figure 4a we plot the steady-state firing rates for all 64 neurons for both networks, while regularly sampling $x_{1}$ and $x_{2}$. RNNs representations are sparse, with only $\sim 10\,\%$ of neurons active at any time, which is in line with sparse coding in the brain. In addition, the average firing rate is $\sim 1$ spike/s, which is surprisingly close to cortical values. Transformers on the other hand, do not have these features, shared by RNNs and their biological counterparts. Furthermore, we find that both networks display mixed selectivity, i.e. neurons are tuned to both variables, which is a known property of cortical neurons [63] (Figure 4b). This suggests that metrics of disentanglement that assume that individual neurons encode distinct factors of variation [30, 38, 13, 32, 21] might be insufficient in detecting disentanglement in networks that generalize well. While recent work incorporates such axis-alignment in the definition of disentaglement, our work along with others [37] showcases the advantages of approaching disentanglement from a mixed representations perspective. Importantly, these properties were not imposed during training, nor was there any parameter fine tuning involved; they emerged from task and optimization objectives.
5.3 Experiments confirm and extend theoretical predictions
@@
Here we expore the relation between the theory in Section 3 and Appendix B and experiments in Section 5 in more depth. First, we wondered why performance saturates in our networks to a high yet non-1 $r^{2}$. The central limit theorem predicts that the estimate of the ground truth $\mathbf{x}^{*}$ in any optimal multi-task classifier becomes more accurate with $\sqrt{t}$, providing a theoretical maximum $r^{2}$ given trial duration $T$ (Section A.4). Since the RNNs trained on the free reaction-time (free RT) task in Section A.10 are required to output their decision confidence at any time in the trial, we can compute OOD $r^{2}$ of free RT network predictions at any timepoint $t$, and compare that to the theoretical prediction. Figure 5a shows that indeed the highest RNN $r^{2}$ falls in the vicinity of or just short of the theoretical maximum. This indicates that RNNs trained with BPTT on these tasks behave like near-optimal multi-task classifiers that create increasingly accurate predictions with time, tightening the relation between our theoretical and experimental results.
@@
Figure 5: Experiments confirm theoretical predictions. (a) OOD $r^{2}$ for free RT RNN required to report its estimate of $\mathbf{x}^{*}$ at different times (see Section A.10 for details, $N_{\textit{task}}=24$, $D=2$). Maximum network $r^{2}$ matches optimal multi-task classifier theory predictions (Equation 8 in Section A.4). (b) OOD $r^{2}$ as a function of input dimensionality $D$ and number of tasks $N_{\textit{task}}$. Good values of $r^{2}$ are obtained when $N_{\textit{task}}\geq D$, especially for GPT models, confirming our theoretical results. (c) Increasing amounts of noise in pretraining results in better OOD generalization ($D=2$).
An important prediction of our theory is that to learn abstract representations, $N_{\textit{task}}$ should exceed $D$. To test this, we increase $D$ (adding more inputs to Figure 2b), while varying $N_{\textit{task}}$. Sampling classification hyperplanes homogeneously (similar to Figure 2a, center) in high-dimensional spaces is non-trivial; therefore we resort to randomly sampling them. Figure 5b shows OOD generalization performance for various combinations of $D$ and $N_{\textit{task}}$. We observe that performance is bad when the $N_{\textit{task}}<D$, but it increases when $N_{\textit{task}}\geq D$. For RNNs, this increase is abrupt for smaller $D$ and more gradual for higher, which is in line with remarks by Johnston and Fusi [37] that it is easier to learn abstract representations when $D$ is high. Transformers on the other hand display higher generalization performance than RNNs, and always perform almost perfectly when $N_{\textit{task}}\geq D$, demonstrating their superior performance in learning abstract representations. Looking at the angles between latents for higher $D$ (Figure S10), we find that transformers have excellent disentanglement as long as $N_{\textit{task}}\geq D$, which might explain their superior generalization performance to RNNs for lower $N_{\textit{task}}$. These results, together with Figure 3c demonstrate the superior ability of transformers in disentangling latent factors. Overall, our findings confirm our theory that abstract representations emerge when $N_{\textit{task}}\geq D$, and even go beyond to suggest that disentangled representations emerge earlier than the theoretical condition $N_{\textit{task}}\gg D$, as long as the architecture is appropriate. These results are remarkable, especially for high $D$, because they go against our intuition that $N_{\textit{task}}$ should scale exponentially with $D$ to fill up the space adequately; instead it need only scale linearly.
Importance of noise for generalization
@@
Our theory and experiments provide insight on the importance of noise for developing efficient, abstract representations (Figure S9e). The closer to a classification boundary the ground truth $\mathbf{x}^{*}$ is, the more likely noise will cross over the boundary. Since, as our theory shows, any optimal multi-task classifier has to estimate $\Pr\{y_{i}(\mathbf{x}^{*})=1\}$, and said probability directly relates to the actual distance from the boundary, it follows that noise allows the model to learn distances from boundaries (Lemma B.3)), leading to efficient localization. We reasoned that additional noise might be even more beneficial, as it would allow more accurate estimation of $\Pr\{y_{i}(\mathbf{x}^{*})=1\}$, especially when $\mathbf{x}^{*}$ is far from the boundary. To test this, we increase noise strength $\sigma$ when pretraining RNNs, while testing with the same $\sigma=0.2$. Indeed, increasing amounts of noise consistently result in better OOD generalization (Figure 5c). This benefit comes for smaller numbers of tasks, allowing us to consider less supervised tasks (e.g. 3), train on them with more noise, and achieve the same performance as more tasks (e.g. 12). So even though networks with more noise perform worse in pretraining (low 90%s classification accuracy), they learn more abstract representations. These findings are highly non-trivial, and have informed our thinking about generalization and inherent variability of the underlying latent factors.